# Personalizing story generation by inferring author styles

### v1

@@ Abstract
Personalization has become essential for improving user experience in interactive writing and educational applications, yet its potential in story generation remains largely unexplored. In this work, we propose a novel two-stage pipeline for personalized story generation. Our approach first infers an author’s implicit story-writing characteristics from their past work and organizes them into an Author Writing Sheet that is inspired by narrative theory. The second stage uses this sheet to simulate the author’s persona through tailored persona descriptions and personalized story writing rules. To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors across 5 distinct sources that reflect diverse story-writing settings. A head-to-head comparison with a non-personalized baseline demonstrates our pipeline’s effectiveness in generating high-quality personalized stories. Our personalized stories achieve a 75% win rate (versus 14% for the baseline and 11% ties) in capturing authors’ writing style based on their past works. Human evaluation highlights the high quality of our Author Writing Sheet and provides valuable insights into the personalized story generation task. Notable takeaways are that writings from certain sources, such as Reddit, are easier to personalize than others, like AO3, while narrative aspects, like Creativity and Language Use, are easier to personalize than others, like Plot.
@@ 1 Introduction
Personalizing the outputs of large language models (LLMs) has promising applications for role-playing in chatbots and video games (Tseng et al., 2024; Chen et al., 2024), as well as providing personalized feedback in education (Stahl et al., 2024; Li et al., 2024b; Makridis et al., 2024). However, story-writing personalization, a crucial writing system feature that can assist users in overcoming writer’s block or help second-language learners adapt to a new language (Yuan et al., 2022; Yeh et al., 2024), remains largely unexplored.
To address this gap, we systematically explore how to personalize LLM outputs based on individual authors’ previous writings. We propose a novel two-stage pipeline for personalized story generation, as shown in Figure 1. In the first stage, we construct an Author Writing Sheet to capture implicit writing characteristics from an author’s past writings (Sun et al., 2025, 2024; Ramos et al., 2024). We categorize story-writing characteristics into four fine-grained narrative dimensions: Plot, Creativity, Development, and Language Use (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024). Each characteristic is represented as a Claim-Evidence pair, where general statements about the author’s writing behavior are supported by specific story excerpts. This approach, inspired by Common Core Standards in English Language Arts (Association et al., 2010) and recent work in writing education (Li et al., 2023), provides a structured and grounded way to identify an author’s distinctive writing patterns, as validated via our human evaluation (Section 3). In the second stage, we implement Personalized Story Generation by having LLMs role-play authors based on their Writing Sheet. Our method uses persona prompts and story rules as user constraints to guide the generation.
@@
To enable and validate our approach, we construct Mythos, a dataset of 590 stories from 64 authors ($\approx$1500 tokens) across five sources with distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker. The dataset splits each author’s works chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
@@
We evaluate personalized stories in two ways: First, we use automatic evaluation with LLM-as-a-judge, to assess faithfulness to the author’s writing history and similarity to ground truth author writings, along with traditional metrics for lexical overlap, story diversity, and style similarity. Second, we conduct a human evaluation to assess the generated story’s similarity to the author’s ground-truth writings. We find that stories generated using Author Writing Sheets demonstrate clear advantages over a non-personalized baseline approach. When evaluated by LLMs, personalized stories have 61% higher faithfulness to the author writing history. Human annotators prefer these stories by an 18% absolute margin in terms of similarity to ground truth. Our personalization approach is especially effective for Reddit content and for capturing Creativity, and Language Use characteristics, with less pronounced impact on Plot (Tian et al., 2024; Xu et al., 2024). In summary, our contributions are:
@@
  • Mythos: A novel dataset containing multiple stories written by the same authors, designed to support both author profiling and personalized story generation evaluation.
@@
  • Author Writing Sheet: A structured approach to capturing an author’s writing behavior based on past works, grounded in narrative theory and educational practices and validated through human evaluation.
  • Personalized Story Generation: A method for generating stories in an author’s style using Author Writing Sheets, rigorously evaluated through automated and human assessments, demonstrating its superiority over baselines.
@@ 2 Dataset Collection and Description
To facilitate our analysis of personalized story generation, we ensemble Mythos, a novel dataset consisting of 590 stories written by 64 authors. To the best of our knowledge, Mythos is the first dataset that identifies and connects multiple stories written by the same author. See Table 1 for comparison with existing datasets.
@@
We include five diverse story-writing sources, namely (1) Reddit, featuring stories from r/WritingPrompts, a widely-used resource for amateur story-writing research (Fan et al., 2018), (2) AO3, featuring fanfiction based on popular franchises such as Harry Potter and Sherlock Holmes, (3) Storium, featuring collaborative stories (Akoury et al., 2020), (4) N.Magazine, featuring professionally polished stories by authors such as Barry Gifford, and finally (5) New Yorker, featuring expert-level storytelling from renowned authors such as Haruki Murakami.
@@
To ensure high-quality and diverse content, we impose constraints on word length (ranging from 500 to 1500 words), limit publication dates to post-November 2023, and exclude explicit material. To standardize formatting, non-Reddit stories are supplemented with writing prompts generated by GPT-4o and manually reviewed for accuracy. Stories are split chronologically based on their submission timestamps, with 70% used for profiling writing characteristics and 30% for generation (see Appendix A for details).
Dataset size:
Our dataset is limited in size due to data crawling policies that restrict large-scale data collection for analysis [^1] https://archiveofourown.org/admin_posts/25888. We note that recent storywriting datasets such as TELL ME A STORY (Huot et al., 2024) and TTCW (Chakrabarty et al., 2024) have conducted analyses on a comparable number of stories (Table 1).
Table 1: Comparison of our dataset with existing story-writing datasets. ✗ indicates no Author IDs associated with stories, and $\sim$ denotes having Author IDs but not having explicit links between them. Our dataset spans diverse settings and links stories by the same author, enabling research on personalized story generation.
@@
TTCW (Chakrabarty et al., 2024) 12 New Yorker 54 2000 $\sim$
Mythos 590 Reddit, AO3, Storium, N.Magazine, New Yorker 50 1517
@@ 3 Author Writing Sheet
Motivated by the connection between writing education and personalization (Li et al., 2023), we propose a novel method to infer implicit story-writing characteristics of authors based on their history in the profiling set $P$. We group these characteristics into four narrative categories: Plot, Creativity, Development, and Language Use, inspired by narrative theory (Pavis, 1998; Card, 1999; Noble, 1994; Huot et al., 2024), to assess fine-grained, multi-dimensional story-writing attributes:
@@
Inspired by Common Core (CC) Standards in English Language and Arts (Association et al., 2010), each characteristic is represented as a Claim-Evidence pair, where the Claim summarizes the author’s writing style, and Evidence consists of excerpts from the author-written story to justify the Claim. These characteristics are organized into an Author Writing Sheet (Gurung and Lapata, 2024).
Stage 1 in Figure 1 (and Algorithm 1 in the Appendix) illustrates our method for generating the Author Writing Sheet by iteratively processing each story in an author’s profiling set. For a given writing prompt ($wp_{t}$) and corresponding author-written story ($s_{at}$), we use an LLM (LLMavg) to generate an average story ($s_{bt}$). Since LLMs are pre-trained on large-scale text data from the web, this average story simulates a typical author’s response based on source-specific stylistic patterns (see Average Author). We prompt an LLM (LLMsheet) to compare the average and author-written stories and generate Claim-Evidence pairs for each narrative category to form an intermediate Author Writing Sheet $A_{t}^{\prime}$, which captures how the author’s writing characteristics deviate from those of a typical author for that source (Shashidhar et al., 2024; Krishna et al., 2020).
To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate Author Writing Sheet $A_{t}^{\prime}$ with the aggregated sheet from prior stories $A_{t-1}$, maintaining a moving update similar to (Chang et al., 2023). During merging, we prompt an LLM (LLMcombine) to integrate $A_{t}^{\prime}$ into $A_{t-1}$ by grouping equivalent Claims within each narrative category, selecting the best Evidence for grouped Claims, and retaining ungrouped Claims with their respective Evidence. To improve clarity, the updated sheet includes a timestamp identifier for each Evidence corresponding to the story it belongs to and also limits each narrative category to 10 Claim-Evidence pairs, prioritizing grouped Claims. For LLMavg, LLMsheet, and LLMcombine we use GPT-4o prompted in a chain-of-thought manner with a temperature of 0.0 and a maximum output of 4096 tokens (Wei et al., 2022; Shashidhar et al., 2024).
Our method processes each story separately, making the update process efficient when new submissions are added (Yeh et al., 2024; Yuan et al., 2022). This method avoids the need to reprocess all previous stories, reducing computational cost and mitigating the limitations of long-context inputs in LLMs (Zhou et al., 2023; Magar and Schwartz, 2022; Li et al., 2024a). This approach draws parallels to Knowledge Tracing in education research, where a model continuously updates its understanding of a student’s progress without re-evaluating past interactions (Liu et al., 2022; Scarlatos and Lan, 2024). See Appendix B.1 for cost analysis.
Validating Author Writing Sheets:
@@
We conduct a human evaluation with three Upwork annotators (\$17/hour). Annotators assess the Claim-Evidence pairs based on two criteria: (1) whether the Claims can be reasonably inferred from the story (Yes/No) and (2) whether the Evidence supports the Claims (Yes/Partially/No). The study includes 42 authors spanning all five sources, each represented by one sampled story resulting in 188 annotated claims. Of these, 12 stories are annotated by all annotators to measure inter-rater agreement, while each annotator evaluates 10 additional exclusive stories. Results show unanimous agreement on the ‘Yes’ label for Claim inference and moderate agreement on Evidence support (Krippendorff’s Alpha=.$=.57$), with 93% of Evidence fully supporting Claims and 7% providing partial support. These findings confirm the reliability and quality of the Author Writing Sheets and support their use in personalized story generation (see Appendix B.2 for more details).
@@ 4 Personalized Story Generation
Stage 2 of Figure 1 illustrates our proposed method for personalized story generation using the Author Writing Sheet.
@@
To generate personalized stories, we prompt an LLM (LLMstory) with the writing prompt, story length (Chakrabarty et al., 2024), source-specific metadata (e.g., fanfiction for AO3), story rules as actionable instructions in direct second-person form (categorized into four narrative categories), and persona description obtained from the Author Writing Sheet. We experiment with four types of methods for personalized story generation: 1) A non-personalized baseline, Average Author; 2) Personalization baselines, RAG and Delta, that do not use the Author Writing Sheet; 3) Our proposed methods, Writing Sheet and Writing Summary; and 4) An Oracle method that loosely resembles the upper bound performance. We detail each method type below. See Appendix C for the prompts.
Non-Personalized - Average Author:
@@
The Average Author method serves as a non-personalized baseline that reflects the average behavior of authors, acquired during an LLM’s pre-training process. Specifically, for each given source, we prompt an LLM to generate stories using an Average Author prompt that describes typical writing characteristics for the given story source obtained using audited GPT-4o prompting [^2] We ask GPT-4o for the typical writing characteristics for the story source and manually verify it. (Wang et al., 2024). For example, for AO3, the Average Author is defined as “respecting fandom tone and style while experimenting with established tropes, unconventional pairings, or alternative universes.”
Personalization Baselines
RAG:
@@
The Retrieval-Augmented Generation, RAG baseline (Salemi et al., 2024) first retrieves the most similar writing prompt and author-written story from the profiling set using BM25 (Robertson et al., 2009). We then use the retrieved pair as a one-shot demonstration (Wang et al., 2024) to elicit role-playing behavior from the LLM, mimicking the retrieved example’s style.
Delta:
@@
The Delta method generates personalized story rules by contrasting the Average Author story (see Average Author) with the corresponding author-written story for each writing prompt in the profiling set (Shashidhar et al., 2024). We then use all writing prompts in the profiling set, along with their corresponding generated story rules that are actionable instructions in direct second-person form, as few-shot demonstrations for the LLM, to generate personalized story rules for a new prompt in the generation set.
Proposed Method - Writing Sheet and Summary
Writing Sheet:
@@
Our Writing Sheet method uses the Author Writing Sheet (Section 3) to enable personalization (Stage 2 of Figure 1). First, we prompt an LLM (LLMpersona) to generate a persona description that summarizes the author’s story-writing style as a second-person narrative, included in the system prompt (Wang et al., 2024; Jiang et al., 2024). Second, we prompt an LLM (LLMrule) to generate personalized story rules from the Author Writing Sheet tailored to the writing prompt in the generation set, included as constraints in the user prompt (Pham et al., 2024). Additionally, we include a one-shot demonstration, following the approach described in RAG (Richardson et al., 2023).
Writing Summary:
@@
As an alternative to the Author Writing Sheet, the Writing Summary method leverages the LLM’s long-context capabilities (Ding et al., 2023) by packing all past stories from the profiling set as input. Specifically, we include them in a single input prompt to generate an Author Writing Summary in the same format as the Author Writing Sheet. Similar to the Writing Sheet method, persona descriptions, and story rules are derived from the Author Writing Summary and used as constraints for story generation, along with a one-shot demonstration.
Oracle:
@@
The Oracle method establishes an upper bound on personalization performance, in terms of similarity to the ground truth, by using Oracle story rules derived directly from the ground-truth author-written story for each writing prompt in the generation set. These rules are obtained by contrasting the Average Author with the ground-truth author story, following the same process as in Delta. Additionally, we include a one-shot demonstration, following the approach described in RAG.

### v2

@@ Abstract
Personalization is critical for improving user experience in interactive writing and educational applications, yet remains understudied in story generation. We study the task of personalizing story generation, where our goal is to mimic an author’s writing style, given other stories written by them. We collect Mythos, a dataset of 3.6k stories from 112 authors, with an average of 16 stories per author, across five distinct sources reflecting diverse story-writing settings. We propose a two-stage pipeline for personalized story generation: first, we infer authors’ implicit writing characteristics and organize them into an Author Writing Sheet, which is validated by humans to be of high quality; second, we simulate the author’s persona using tailored persona descriptions and personalized story rules. We find that stories personalized using the Author Writing Sheet outperform a non-personalized baseline, achieving a 78% win-rate in capturing authors’ past style and 59% in similarity to ground-truth author stories. Human evaluation supports these findings and further highlights trends, such as Reddit stories being easier to personalize, and the Creativity and Language Use aspects of stories being easier to personalize than the Plot.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used in interactive systems to support human users in co-writing tasks (Fang et al., 2024; Yuan et al., 2022; Chung et al., 2022). A key challenge in these settings to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025). Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023). Despite these potentials, personalization in story writing has received limited attention.
@@
Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization, the inherent subjectivity in user preferences, and methods to evaluate the extent of personalization. To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker. We split each author’s stories chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
We systematically explore personalized story generation using LLMs with a two-stage pipeline as shown in Figure 1: First, we construct an Author Writing Sheet to extract implicit writing characteristics from an author’s past stories (Ramos et al., 2024). The Author Writing Sheet covers four narrative dimensions: Plot, Creativity, Development, and Language Use (Pavis, 1998; Huot et al., 2024), and is organized as Claim-Evidence pairs grounded in story excerpts to summarize author preferences. This representation, inspired by Common Core Standards (Association et al., 2010) and writing education research (Li et al., 2023), offers a cost-effective and interpretable way to model authors’ writing preferences. Second, we guide LLMs to role-play authors using persona prompts and story rules derived from the Author Writing Sheet.
@@
We evaluate generated personalized stories in two ways: First, we use automatic evaluation with LLM-as-a-judge to assess faithfulness to the author’s writing history and similarity to ground truth author story, along with traditional text similarity metrics on lexical overlap, story diversity, and style similarity. Second, we conduct a human evaluation to assess the generated story’s similarity to the author’s actual writings.
We find that stories generated using the Author Writing Sheet outperform a non-personalized baseline: LLMs prefer them 78% of the time in faithfulness and 59% in similarity to ground truth, while human annotators prefer them 56% of the times for similarity to ground truth. Our method is especially effective for Reddit stories and in capturing Creativity and Language Use, with a smaller impact on Plot. In summary, our contributions are:
@@
  • Mythos: A dataset of 3.6k stories with multiple stories per author, supporting both author profiling and personalized story generation evaluation.
@@
  • Author Writing Sheet: A cost-efficient, structured method for capturing author characteristics, grounded in narrative theory and validated through human evaluation.
  • Personalization Evaluation: An evaluation combining automated and human assessments, showing the benefits of Author Writing Sheets.
@@ 2 Dataset Collection and Description
To facilitate our analysis of personalized story generation, we assemble Mythos, a dataset consisting of 3.6k stories written by 112 authors. To the best of our knowledge, Mythos is the first dataset that identifies and connects multiple stories written by the same author (see Table 1 for comparison).
@@
We include five diverse story-writing sources: (1) Reddit, featuring stories from r/WritingPrompts, a widely-used resource for amateur story-writing research (Fan et al., 2018), forming the largest subset with 3.2k stories, (2) AO3, featuring fanfiction based on popular franchises like Harry Potter, (3) Storium, featuring collaborative stories (Akoury et al., 2020), (4) N.Mag (Narrative Magazine), featuring professional stories from authors like Barry Gifford, and (5) N.York (New Yorker), featuring expert-level storytelling from authors like Haruki Murakami.
@@
To ensure high-quality and diverse content, we impose constraints on story length (ranging from 500 to 1500 words), limit publication dates to post-November 2022, and exclude stories with explicit content. To standardize formatting, non-Reddit stories are supplemented with writing prompts generated by GPT-4o, which are manually reviewed for accuracy. Stories are split chronologically, with 70% used for profiling writing characteristics and 30% for testing generation. Details on our dataset can be found in Appendix A.
@@
Mythos (ours) 3.6k Reddit, AO3, Storium, N.Mag, N.York 50 1517
Table 1: Comparison of our dataset with existing story-writing datasets. ✗ indicates no Author IDs associated with stories, and $\sim$ denotes having Author IDs but not having explicit links between them. Mythos covers diverse story-writing settings and links stories by the same author, enabling future research on personalized story generation.
@@ 3 Author Writing Sheet
Motivated by the connection between writing education and personalization (Li et al., 2023), we propose a method to infer authors’ implicit story-writing tendencies based on their past stories in the profiling set, denoted as $P$. We group these tendencies into four narrative categories: Plot, Creativity, Development, and Language Use inspired by narrative theory (Pavis, 1998; Huot et al., 2024), to capture fine-grained, multi-dimensional aspects of how authors approach story-writing:
@@
Inspired by the Common Core Standards in English Language Arts (Association et al., 2010), each characteristic is represented as a Claim-Evidence pair, where the Claim summarizes the author’s style and the Evidence are supporting story excerpts.
Stage 1 in Figure 1 (and Algorithm 1 in the Appendix) illustrates our method for generating the Author Writing Sheet (example in Table 7), where we iteratively process each story in an author’s profiling set.
We construct the sheet by merging characteristics across stories that highlight how the author’s style differs from that of a typical, average author. For each writing prompt ($wp_{t}$) and its corresponding author story ($s_{at}$), we use the LLM with two distinct prompts: one prompt (LLMavg) elicits an average story ($s_{bt}$) that reflects the typical style of an author on that writing source, relying solely on source-specific priors acquired during pre-training; another prompt (LLMsheet) asks the same LLM to compare the author story ($s_{at}$) with the average story ($s_{bt}$) and extract corresponding Claim-Evidence pairs across narrative categories. This yields an intermediate Author Writing Sheet $A_{t}^{\prime}$ (Shashidhar et al., 2024; Krishna et al., 2020).
To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2023). To reduce redundancy and manage output length, we prompt the LLM with a prompt (LLMcombine) to group equivalent Claims, select the best Evidence, and retain ungrouped Claims. We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations. Each Evidence is tagged with a timestamp indicating its source story.
Similar to Knowledge Tracing in education (Liu et al., 2022; Scarlatos and Lan, 2024), our method processes each story sequentially. Doing so enables efficient updates to the Author Writing Sheet without having to re-process prior stories every time a new story is analyzed (Yeh et al., 2024). Therefore, this method reduces cost and mitigates long-context limitations in LLMs (Li et al., 2024); see Appendix B.1 for a detailed cost analysis.
Validating Author Writing Sheet Quality:
@@
We conduct a human evaluation with three Upwork annotators to assess the quality of the Claim-Evidence pairs. Annotators evaluate each pair on two criteria: (1) whether the Claim can be reasonably inferred from the story (Yes/No), and (2) whether the Evidence supports the Claim (Yes/Partially/No). The study spans 42 authors across five sources, with one sampled story per author, yielding 188 annotated claims. To measure inter-rater reliability, 12 stories are annotated by all annotators. Results show unanimous agreement on the ‘Yes’ label of Claims and moderate agreement on Evidence support (Krippendorff’s $\alpha=0.57$), despite the task’s subjectivity. Overall, 93% of Evidence statements fully support their Claims, and the remaining 7% provide partial support. These results indicate high precision in constructing Claim-Evidence pairs and confirm the reliability and quality of the Author Writing Sheet; see Appendix B.2 for more details including experimental design and interface (Figure 13), pilot studies, and results.
@@ 4 Personalized Story Generation
We now detail our proposed method for personalized story generation using the Author Writing Sheet, illustrated in Stage 2 of Figure 1. To generate personalized stories, we prompt an LLM (LLMstory) with the writing prompt, story length, source-specific metadata (e.g., fanfiction for AO3), story rules as actionable instructions in direct second-person form categorized into four narrative categories (example in Table 8), and persona descriptions obtained from the Author Writing Sheet (example in Table 9). We experiment with four types of methods: (1) A non-personalized baseline, Average Author; (2) Personalization baselines, RAG and Delta, that do not use the Author Writing Sheet; and (3) Our proposed methods, Sheet and its variant Summary. (4) An Oracle method that loosely resembles the upper bound performance; See Appendix C for the prompts used.
Non-Personalized: Average Author
@@
The Average Author method serves as a non-personalized baseline reflecting typical author behavior learned during LLM pre-training. For each source, we prompt an LLM with an Average Author description summarizing writing characteristics obtained through audited GPT-4o prompting, [^1] We ask GPT-4o for typical writing characteristics for each source and manually verify them. verified manually (Wang et al., 2024). Example, for AO3, the Average Author is defined as “respecting fandom tone and style while experimenting with tropes, unconventional pairings, and alternate universes.”
Personalization Baselines: RAG and Delta
RAG
@@
The Retrieval-Augmented Generation (RAG) baseline (Salemi et al., 2024) first retrieves the most similar writing prompt and author story from the profiling set using BM25 (Robertson et al., 2009). Following (Wang et al., 2024), the retrieved pair is used as a one-shot demonstration to elicit role-playing behavior from the LLM, mimicking the retrieved example’s style.
Delta
@@
The Delta method generates personalized story rules by contrasting the Average Author story (see Average Author) with the corresponding author-written story for each writing prompt in the profiling set. Following (Shashidhar et al., 2024), we then use all writing prompts in the profiling set, along with their corresponding generated story rules that are actionable instructions in direct second-person form, as few-shot demonstrations for the LLM, to generate personalized story rules for a new prompt in the generation set.
Our Methods: Sheet and Summary
Sheet
@@
Our Sheet method uses the Author Writing Sheet (Section 3) for personalization (Stage 2 of Figure 1). First, we prompt an LLM (LLMpersona) to generate a persona description that summarizes the author’s story-writing style as a second-person narrative, included in the system prompt (Wang et al., 2024; Jiang et al., 2024). Second, we prompt an LLM (LLMrule) to generate personalized story rules from the Author Writing Sheet tailored to the writing prompt in the generation set, included as constraints in the user prompt (Pham et al., 2024) for LLMstory. Additionally, following (Richardson et al., 2023), we include a one-shot demonstration using the approach described in RAG.
Summary (Summ)
@@
As a variant of the Author Writing Sheet, the Summ method leverages the LLM’s long-context capabilities (Ding et al., 2023) by packing all past stories from the profiling set as input in the same prompt to generate an Author Writing Summary in the same format as the Author Writing Sheet. Similar to the Sheet method, persona descriptions and story rules are derived from the Author Writing Summary and used as constraints for story generation, along with a one-shot demonstration.
Reference Method: Oracle
@@
The Oracle method establishes an upper bound on personalization performance by using story rules derived directly from the ground-truth author story for each writing prompt. These rules are obtained by contrasting the Average Author with the ground-truth story (see Delta). We also include a one-shot demonstration following the approach described in RAG.